# StoneDetector: Conventional and versatile code clone detection for Java

### v1

@@ 4.5 Research Question 4: Scalability
Scalability is another important requirement for clone detectors, i.e., such tools need to be able to process large quantities of code in reasonable time. In order to evaluate StoneDetector’s capabilities of analyzing large fractions of code, we use a setup described by Sajnani et al. [18]. To this end, random subsets of files from the IJaDataset 2.0 [45] are generated, comprising 1, 10, and 100 million lines of code, respectively. We ran StoneDetector and other clone detectors on these subsets and report their runtime in Table 6. As can be seen, Deckard has been omitted, since Deckard’s memory consumption exceeds available resources, as has also been noted by Sajnani et al. [18]. Furthermore, we observe that CCAligner as well as iClones are not able to scale to 100 million lines of code (cf. also [10]). Also, contradicting previous results in [18], the clone detector NiCad in its recent version NiCad v6.2 scales to more than 100 million lines of code and also provides the lowest runtime. We also see that StoneDetector keeps up well with the other state-of-the-art tools and analyzes 100 million lines of code in 7 hours and 13 minutes and 10 million lines of code in under 7 minutes. Using a different string metric as in the default configuration allows StoneDetector to run even faster but also partially sacrifices recall and precision (cf. Sect. 4.7).
Due to these encouraging results, we performed another scalability check using StoneDetector on the entire IJaDataset 2.0 [45], comprising approx. 2.5 million Java source files with 328 million lines of code. StoneDetector has been able to analyze this very large dataset in 79 hours. Again note that we report on an even faster version of StoneDetector when the Hamming distance is used as a string metric, as described in Sect. 4.7.
StoneDetector scales to large codebases comprising multiple 100 million lines of code and features runtimes comparable to the state of the art.
4.6 Research Question 5: Configuration
SourcererCC NiCad CCAligner StoneDetector NIL Oreo
Recall 35,843 35,847 35,744 35,843 35,783 35,480
T 1 99.9% 99.9% 99.6% 99.9% 99.8% 98.9%
Recall 4,432 4,550 4,530 4,579 4,421 4,506
T 2 96.7% 99.3% 98.8% 99.9% 96.4% 98.3%
Recall 3,814 4,098 4,055 4,140 3,892 3,893
VST 3 91.6% 99.2% 97.4% 99.4% 93.5% 93.5%
Recall 9,026 10,680 10,534 13,008 10,083 12,625
ST 3 59.2% 70.0% 69.1% 85.3% 66.1% 82.8%
Recall 3,935 476 8,099 14,537 8,548 22,965
MT 3 4.8% 0.6% 10.0% 17.9% 10.5% 28.2%
Recall 1,690 11 12,673 11,612 14,500 52,839
T 4 0.0% 0.0% 0.2% 0.1% 0.2% 0.7%
@@ 4.6 Research Question 5: Configuration
Clone detection using StoneDetector should result in the highest possible recall with acceptable precision and runtime. In order to achieve this objective, certain parameters of our clone detection method have to be chosen. Thus, our fifth research question addresses experiments on optimal configuration parameters in StoneDetector, in particular considering the threshold values for syntactical similarity of code clones and their minimum size.
In the experiments presented so far, we have chosen 15 lines as the minimum size of code clone fragments. In order to justify our previous claim that smaller code fragments in general result in a skewed comparison, we performed additional experiments using different minimum sizes of code clones on BigCloneBench. In Table 7, we report recall values when a minimum size of 6 lines of code is used instead of 15 lines, since 6 lines of code are often used in the literature as well. Note that we have again omitted the tools Deckard and iClones due to their lower precision. As expected, the shown recall values are in general higher compared to those in Table 4. This can be well explained by Java methods with only 6 lines of source code, where half of the lines are just taken by the method’s annotations, signature, and curly braces. We, though, do not observe substantial changes in the tools’ comparisons, besides Oreo now outperforming StoneDetector in the case of Moderately Type 3 code clones. Furthermore, we also considered other minimum sizes in the range from 6 to 20 lines of code, though only for our tool StoneDetector. Due to space constraints, instead of providing the respective numbers, we have plotted the resulting precision and recall on BigCloneBench in Fig. 6. As can be seen, comparable recall values have been obtained, while precision values range from 96% to 100%, with increasing values for increasing minimum sizes of code clones. We can thus confirm our stated conclusion that increased minimum clone size supports the precision of clone detectors [10].
@@
A second parameter we are interested in is the similarity threshold, which is used in our method to define whether two processed code fragments are considered code clones or not (cf. also Sect. 2.5). In the experiments conducted so far, this threshold value has been set to 70%, i.e., distance value $\tau=0.3$, which on the one-hand side is comparable to the threshold values as used in other clone detectors and on the other-hand side, in our experience, provides an optimal tradeoff of high recall and precision. In order to confirm this optimal tradeoff, we conduct additional experiments using StoneDetector on BigCloneBench using differing threshold values in terms of the distance value $\tau$ ranging from $\tau=0.1$ to $\tau=0.4$. The respective distance values are again plotted in conjunction with the resulting recall and precision in Fig. 6 (other configuration parameters equal those in Table 2, minimal size of code clones 15 lines of code). As expected, if the distance value increases, the recall increases with a simultaneous decrease in precision. The precision values range between 96% and 100%. It is striking that with low distance values, the recall values for Type 3 and Type 4 code clones are also lower. In these cases, the distance value is lower than the admitted syntactical difference of code clones. However, with increasing distance values, more code clones of Type 3 and Type 4 can be found again, though with a higher inaccuracy. As these results show, we can observe that recall and precision are balanced with a distance value of $\tau=0.3$, i.e., a similarity threshold of $70\%$.
@@
The abstraction of paths in StoneDetector, e.g., the uniform treatment of variables and constants as discussed in Sect. 2, can lead to several identical paths in one description set. As an example, consider two paths that only differ from each other in their last node, and where the respective nodes each represent assignments with which the same/different variables are assigned to different/the same constants. By abstracting variables and constants, these assignments are then mapped to the same representation, which results in an identical abstraction for both paths. We observed such presence of identical paths in one and the same description set to occur quite frequently. Although such occurrences do not influence StoneDetector’s recall or precision in finding code clones, its runtime is deteriorated due to the redundant comparisons of identical paths. Therefore, in additional experiments, we compare a variant of StoneDetector that does not handle identical paths and a variant where two identical paths or two paths that differ only in one node are merged into a single path. Surprisingly, this rather simple strategy already leads to a relative runtime improvement of 20%, which corresponds to a reduction from approx. 25 to 20 minutes on the whole BigCloneBench.
Clone detection on source code with StoneDetector achieves the best results using a similarity threshold of 70% ($\tau=0.3$). Increasing the minimal size of clones in general increases the precision of clone detectors.
4.7 Research Question 6: Metrics and Hashing
@@
Another parameter that can impact StoneDetector’s capability in finding code clones is the similarity measure, i.e., string metric, used for comparing the paths of two code fragments’ dominator trees (cf. also Sect. 2.3). Furthermore, since the path comparison is working on hashed fingerprints of the paths, the utilized hashing mechanism may as well influence StoneDetector’s performance. For our sixth research question, we therefore want to specifically analyze the role of the string metric and hashing mechanism. In our previously described experiments, we used a combination of a variant of the Longest Common Subsequence (LCS)-based metric ($\delta_{LCS_{modified}}$) in combination with Locality-Sensitive Hashing (LSH). This is in contrast to the authors’ previous work in [21], where we used the standard LCS-based metric in combination with a 4-byte prime number-based hashing mechanism.
@@
In a first experiment, we considered the role of using different string metrics on clone detection in StoneDetector. To this end, we have chosen a subset of the string metrics implemented in StoneDetector, i.e., the LCS-based metric, modified LCS-based metric, Levenshtein distance, Needleman-Wunsch, Hamming distance, and modified Hamming distance. The latter metric depicts a variant of the Hamming distance in which different path lengths are not penalized, in the sense that the difference between different paths is not included in the distance measure as a penalty. We refer the interested reader to Sect. 2.3 for a short explanation of these string metrics. We analyze the role of a certain string metric by evaluating StoneDetector’s recall and precision on BigCloneBench when using the respective metric and MD5 hashing, keeping the other parameters stable (cf. Table 2). The results of this first experiment are shown in Table 8. Note that we did not observe any differences in the results when using either MD5 or a simple 4-byte hashing and are therefore denoting the results for the standard LCS-based metric with 4-byte hashing for the sake of comparison with our previous work [21]. In summary, the shown results indicate that using the Needleman-Wunsch algorithm or the modified LCS-based metric achieves the best performance for clone detection. Compared with these metrics, the use of the Levenshtein distance yields a visible deterioration in the recall values for Moderately and Strongly Type 3 code clones, while precision remains the same. In contrast, the Hamming distance implies significantly poorer recall values when looking at the syntactically more deviating code clones, i.e., clones of Type 3 and Type 4. While the use of the modified Hamming distance results in improved recall values, it features the worst precision. However, we observed significantly shorter runtimes of StoneDetector in the case of using either variant of the Hamming distance. Using the Hamming distance resulted in a speed-up of approximately 3.5 compared to the other metrics. As a rule of thumb to be taken from these measurements, we can thus state that the Hamming distance should be preferred over the other metrics when searching exact and near-miss clones, i.e., Type 1 and Type 2 code clones, due to its impact on shortening StoneDetector’s runtime.
@@
In a second experiment, the hashing algorithm’s influence on StoneDetector’s performance is evaluated. Since the modified LCS-based metric has so far achieved the best results, we consider only this string metric in this experiment and compare the tool’s recall and precision when using either the MD5 hashing algorithm or Locality-Sensitive Hashing (LSH). Note again that we did not observe differences when using MD5 or a simple 4-byte hashing. As can also be seen in Table 8, the usage of LSH results in particular in improved recall values for Moderately Type 3 and Type 4 code clones, while maintaining precision. We have therefore chosen the modified LCS-based metric and LSH as StoneDetector’s default configuration in our evaluation.
@@
Clone detection on source code with StoneDetector achieves the best results using the modified Longest Common Subsequence-based string metric in combination with Locality-Sensitive Hashing. Using the Hamming distance instead, though, allows for faster clone detection.
4.8 Research Question 7: Java Bytecode
@@
Clone detection for Java Bytecode can be done either directly using the raw Bytecode or using an intermediate representation as generated, e.g., during compilation. The Soot framework [41] supports the transformation of Java Bytecode into several intermediate representations. Soot’s Baf format roughly resembles the original Bytecode and provides a stack-based intermediate representation. In contrast, the Jimple format is a register-based three-address code, which is generated by Soot out of the stack-based Bytecode for program analysis and transformation. We consider both formats in our evaluation and are interested in the their different suitability for code clone detection. Evaluating clone detectors for Java Bytecode is unfortunately not as straightforward as for Java source code, because the state-of-the-art benchmark BigCloneBench consists of only plain source files and is missing their respective dependencies [40]. We therefore use the Stubber tooling to transform BigCloneBench’s Java source files in such a way, that they can be compiled to Java Bytecode without the missing dependencies. For the details, we refer the interested reader to the publication introducing Stubber [47]. Using Stubber allows us to generate Bytecode for 52,816 or 95% of the overall 54,999 benchmark’s source files and to subsequently generate the Bytecode’s Baf and Jimple representation using Soot.

### v2

@@ 4.4 Research Question 3: Scalability
Scalability is another important requirement for clone detectors, i.e., such tools need to be able to process large quantities of code in reasonable time. In order to evaluate StoneDetector’s capabilities of analyzing large fractions of code, we use a setup described by Sajnani et al. [88]. To this end, random subsets of files from the IJaDataset 2.0 [7] are generated, comprising 1, 10, and 100 million lines of code, respectively. We ran StoneDetector and other clone detectors on these subsets and report their runtime in Table 6. As can be seen, Deckard has been omitted, since Deckard’s memory consumption exceeds available resources, as has also been noted by Sajnani et al. [88]. Furthermore, we observe that CCAligner as well as iClones are not able to scale to 100 million lines of code (cf. also [110]). Also, contradicting previous results in [88], the clone detector NiCad in its recent version NiCad v6.2 scales to more than 100 million lines of code and also provides the lowest runtime. We also see that StoneDetector keeps up well with the other state-of-the-art tools and analyzes 100 million lines of code in 7 hours and 13 minutes and 10 million lines of code in under 7 minutes. Using a different string metric as in the default configuration allows StoneDetector to run even faster but also partially sacrifices recall and precision (cf. Sect. 4.5).
Due to these encouraging results, we performed another scalability check using StoneDetector and the other clone detectors on the entire IJaDataset 2.0 [7], comprising approx. 2.5 million Java source files with 328 million lines of code. The respective runtime results are as well shown in Fig. 6. As can be observed, only four of the tools have been able to analyze the whole dataset. While CCAligner and iClones already did not scale for 100 million lines of code, NiCad failed for 328 million lines of code due to an internal threshold of input size and SourcererCC was cut off after running for a whole week. Overall, NIL and CloneWorks provide the best runtime, with NIL only requiring less than 2 hours for analyzing the whole IJaDataset. StoneDetector has been able to analyze this very large dataset in 79 hours. Due to its focus on finding harder-to-detect code clones, StoneDetector thus trades higher analysis time for recall. However, note that we report on a faster version of StoneDetector when the Hamming distance is used as a string metric with MD5 hashing, as described in Sect. 4.5, which makes it a fast near-miss clone detector. We have included the respective runtime for this version of StoneDetector in Fig. 6 as well. As can be seen, while not as fast as NIL or CloneWorks, StoneDetector then can analyze 328 million lines of code in 12 and a half hours. Like discussed in Sect. 4.5, this version of StoneDetector though yields significantly poorer recall values.
StoneDetector scales to large codebases comprising multiple 100 million lines of code. Using the Hamming distance instead of the modified LCS metric makes StoneDetector behave like a fast near-miss clone detector.
4.5 Research Question 4: Tool Configuration
SourcererCC NiCad CCAligner StoneDetector NIL CloneWorks Oreo
Recall 35,843 35,847 35,744 35,843 35,783 35,844 35,480
T 1 99.9% 99.9% 99.6% 99.9% 99.8% 99.9% 98.9%
Recall 4,432 4,550 4,530 4,579 4,421 4,550 4,506
T 2 96.7% 99.3% 98.8% 99.9% 96.4% 99.3% 98.3%
Recall 3,814 4,098 4,055 4,140 3,892 4,094 3,893
VST 3 91.6% 99.2% 97.4% 99.4% 93.5% 98.3% 93.5%
Recall 9,026 10,680 10,534 13,008 10,083 14,165 12,625
ST 3 59.2% 70.0% 69.1% 85.3% 66.1% 92.9% 82.8%
Recall 3,935 476 8,099 14,537 8,548 2,782 22,965
MT 3 4.8% 0.6% 10.0% 17.9% 10.5% 3.4% 28.2%
Recall 1,690 11 12,673 11,612 14,500 46 52,839
T 4 0.0% 0.0% 0.2% 0.1% 0.2% 0.0% 0.7%
@@ 4.5 Research Question 4: Tool Configuration
Clone detection using StoneDetector should result in the highest possible recall with acceptable precision and runtime. In order to achieve this objective, certain parameters of our clone detection method have to be chosen. Thus, our fourth research question addresses experiments on optimal configuration parameters in StoneDetector, in particular considering the threshold values for syntactical similarity of code clones, their minimum size, the used similarity measure, i.e., string metric, and hashing algorithm for comparing and encoding code fragments, respectively.
In the experiments presented so far, we have chosen 15 lines as the minimum size of code clone fragments. In order to justify our previous claim that smaller code fragments in general result in a skewed comparison, we performed additional experiments using different minimum sizes of code clones on BigCloneBench. In Table 7, we report recall values when a minimum size of 6 lines of code is used instead of 15 lines, since 6 lines of code are often used in the literature as well. Note that we have again omitted the tools Deckard and iClones due to their lower precision. As expected, the shown recall values are in general higher compared to those in Table 4. This can well be explained by Java methods with only 6 lines of source code, where half of the lines are just taken by the method’s annotations, signature, and curly braces. We, though, do not observe substantial changes in the tools’ comparisons, besides Oreo now outperforming StoneDetector in the case of Moderately Type 3 code clones. Furthermore, we also considered other minimum sizes in the range from 6 to 20 lines of code, though only for our tool StoneDetector. Due to space constraints, instead of providing the respective numbers, we have plotted the resulting precision and recall on BigCloneBench in Fig. 6. Note that we provide the detailed results as supplementary information online [^18] https://stonedetector.fmi.uni-jena.de/. As can be seen, comparable recall values have been obtained, while precision values range from 96% to 100%, with increasing values for increasing minimum sizes of code clones. We can thus confirm our stated conclusion that increased minimum clone size supports the precision of clone detectors [110].
@@
A second parameter we are interested in is the similarity threshold, which is used in our method to define whether two processed code fragments are considered code clones or not (cf. also Sect. 3.5). In the experiments conducted so far, this threshold value has been set to 70%, i.e., distance value $\tau=0.3$, which on the one-hand side is comparable to the threshold values as used in other clone detectors and on the other-hand side, in our experience, provides an optimal tradeoff of high recall and precision. In order to confirm this optimal tradeoff, we conduct additional experiments using StoneDetector on BigCloneBench using differing threshold values in terms of the distance value $\tau$ ranging from $\tau=0.1$ to $\tau=0.4$. The respective distance values are again plotted in conjunction with the resulting recall and precision in Fig. 6 (other configuration parameters equal those in Table 2, minimal size of code clones 15 lines of code). As expected, if the distance value increases, the recall increases with a simultaneous decrease in precision. The precision values range between 96% and 100%. It is striking that with low distance values, the recall values for Type 3 and Type 4 code clones are also lower. In these cases, the distance value is lower than the admitted syntactical difference of code clones. However, with increasing distance values, more code clones of Type 3 and Type 4 can be found again, though with a higher inaccuracy. As these results show, we can observe that recall and precision are balanced with a distance value of $\tau=0.3$, i.e., a similarity threshold of $70\%$.
@@
The abstraction of paths in StoneDetector, e.g., the uniform treatment of variables and constants as discussed in Sect. 3, can lead to several identical paths in one description set. As an example, consider two paths that only differ from each other in their last node, and where the respective nodes each represent assignments with which the same/different variables are assigned to different/the same constants. By abstracting variables and constants, these assignments are then mapped to the same representation, which results in an identical abstraction for both paths. We observed such presence of identical paths in one and the same description set to occur quite frequently. Although such occurrences do not influence StoneDetector’s recall or precision in finding code clones, its runtime is deteriorated due to the redundant comparisons of identical paths. Therefore, in additional experiments, we compare a variant of StoneDetector that does not handle identical paths and a variant where two identical paths or two paths that differ only in one node are merged into a single path. Surprisingly, this rather simple strategy already leads to a relative runtime improvement of 20%, which corresponds to a reduction from approx. 25 to 20 minutes on the whole BigCloneBench.
@@
Another parameter that can impact StoneDetector’s capability in finding code clones is the similarity measure, i.e., string metric, used for comparing the paths of two code fragments’ dominator trees (cf. also Sect. 3.3). Furthermore, since the path comparison is working on hashed fingerprints of the paths, the utilized hashing mechanism may as well influence StoneDetector’s performance. We therefore also want to specifically analyze the role of the string metric and hashing mechanism. In our previously described experiments, we used a combination of a variant of the longest common subsequence-based metric ($\delta_{LCS_{modified}}$) in combination with Locality-Sensitive Hashing (LSH). This is in contrast to the authors’ previous work in [8], where we used the standard LCS-based metric in combination with a simple 4-byte prime number-based hashing mechanism.
@@
In a third experiment, we considered the role of using different string metrics on clone detection in StoneDetector. To this end, we have chosen a subset of the string metrics implemented in StoneDetector, i.e., the LCS-based metric, modified LCS-based metric, Levenshtein distance, Needleman-Wunsch, Hamming distance, and modified Hamming distance. The latter metric depicts a variant of the Hamming distance in which different path lengths are not penalized, in the sense that the difference in length between different paths is not included in the distance measure as a penalty. We refer the interested reader to Sect. 3.3 for a short explanation of these string metrics. We analyze the role of a certain string metric by evaluating StoneDetector’s recall and precision on BigCloneBench when using the respective metric and MD5 hashing, keeping the other parameters stable (cf. Table 2). Precision is again measured by manual validation of a random sample.
The results of this third experiment are shown in Table 8. Note that we did not observe any differences in the results when using either MD5 or a simple 4-byte hashing and are therefore denoting the results for the standard LCS-based metric with 4-byte hashing for the sake of comparison with our previous work [8]. In summary, the shown results indicate that using the Needleman-Wunsch algorithm or the modified LCS-based metric achieves the best performance for clone detection. Compared with these metrics, the use of the Levenshtein distance yields a visible deterioration in the recall values for Moderately and Strongly Type 3 code clones, while precision remains the same. In contrast, the Hamming distance implies significantly poorer recall values when looking at the syntactically more deviating code clones, i.e., clones of Type 3 and Type 4. While the use of the modified Hamming distance results in improved recall values, it features the worst precision. However, we observed significantly shorter runtimes of StoneDetector in the case of using the Hamming distance, as reported in Sect. 4.4 and shown in Table 6. As a rule of thumb to be taken from these measurements, we can thus state that the Hamming distance should be preferred over the other metrics when searching exact and near-miss clones, i.e., Type 1 and Type 2 code clones, due to its impact on shortening StoneDetector’s runtime.
@@
In a fourth experiment, the hashing algorithm’s influence on StoneDetector’s performance is evaluated. Since the modified LCS-based metric has so far achieved the best results, we consider only this string metric in this experiment and compare the tool’s recall and precision when using either the MD5 hashing algorithm or Locality-Sensitive Hashing (LSH). Note again that we did not observe differences when using MD5 or a simple 4-byte hashing. As can also be seen in Table 8, the usage of LSH results in particular in improved recall values for Moderately Type 3 and Type 4 code clones, while maintaining precision. We have therefore chosen the modified LCS-based metric and LSH as StoneDetector’s default configuration in our evaluation.
@@
Clone detection on source code with StoneDetector achieves the best results using a similarity threshold of 70% ($\tau=0.3$). Increasing the minimal size of clones in general increases the precision of clone detectors. StoneDetector performs best using the modified longest common subsequence (LCS)-based string metric in combination with locality-sensitive hashing (LSH). Using the Hamming distance instead, though, allows for faster clone detection.
4.6 Research Question 5: Java Bytecode
@@
Clone detection for Java Bytecode can be done either directly using the raw Bytecode or using an intermediate representation as generated, e.g., during compilation. The Soot framework [107] supports the transformation of Java Bytecode into several intermediate representations. Soot’s Baf format roughly resembles the original Bytecode and provides a stack-based intermediate representation. In contrast, the Jimple format is a register-based three-address code, which is generated by Soot out of the stack-based Bytecode for program analysis and transformation. We consider both formats in our evaluation and are interested in their different suitability for code clone detection. Evaluating clone detectors for Java Bytecode is unfortunately not as straightforward as for Java source code, because the state-of-the-art benchmark BigCloneBench consists of only plain source files and is missing their respective dependencies [94]. We therefore use the Stubber tooling to transform BigCloneBench’s Java source files in such a way, that they can be compiled to Java Bytecode without the missing dependencies. For the details, we refer the interested reader to the publication introducing Stubber [91]. Using Stubber allows us to generate Bytecode for 52,816 or 95% of the overall 54,999 benchmark’s source files and to subsequently generate the Bytecode’s Baf and Jimple representation using Soot.
@@
4.7 Summarization
Our evaluation of StoneDetector in comparison with state-of-the-art conventional clone detection tools illustrates its strengths and its limitations. While near-miss clone detectors like NiCad, SourcererCC, and CloneWorks excel in identifying Type 1 and Type 2 code clones with high precision and fast runtimes (cf. Sect. 4.2, Sect. 4.4), StoneDetector achieves in its default configuration competitive performance results, though not providing for similar fast clone detection runtimes. StoneDetector however demonstrates its capabilites in finding harder-to-detect clones of Type 3 and Type 4 with acceptable precision in reasonable time, also when compared to advanced and/or Deep Learning tools such as NIL and Oreo, as exemplified on benchmarks Google Code Jam and Project CodeNet where StoneDetector achieved the best overall recall (cf. Sect. 4.3). The manual qualitative examination of code clones revealed that the tool in fact finds certain types which are not detected by any of the other tools (cf. Sect 4.2.1). Furthermore, StoneDetector can be adapted to other configurations (cf. Sect. 4.5), e.g., using the Hamming distance metric for scalable and fast near-miss clone detection, which makes it a versatile and rich platform for Java code clone detection and experimentation, including Bytecode-based clone detection (cf. Sect. 4.6).
On the one hand, StoneDetector in its default configuration is thus most beneficial in software maintenance scenarios, where developers are challenged by growing inconsistencies across clone instances [118]. Note that such syntactically more deviating clone instances have also been shown to be prone to delays in propagating code changes, referred to as late propagation [67], potentially resulting in bugs, which makes their detection and tracking even more important. In these cases, StoneDetector’s tolerance for refactored statements, i.e., structural code clones, and larger code edits or subsumed/embedded code fragments, i.e., subclones, enable it to capture clones and therefore bugs that other clone detectors may miss, as revealed in our evaluation. Such scenarios are especially relevant in software maintenance, where inconsistent and insufficient identification of code clones has been shown to correlate with fault proneness [38]. StoneDetector’s ability to detect structurally modified code clones also supports consolidation of functional and architectural adaptations [9], which is less feasible with clone detectors optimized for exact or near-miss clones. On the other hand, StoneDetector allows for other configurations. For example, in certain scenarios, the primary objective of clone detection may be the rapid identification of exact duplicates or near-miss clones in large codebases. Note that such code clones often constitute a large fraction in practice [64]. As mentioned above, using the Hamming distance metric and MD5 hashing makes StoneDetector behave like a fast near-miss clone detector, yielding a speed-up of approx. 6.3, though trading accuracy for speed. Apart from StoneDetectur thus being adaptable to software maintenance scenarios, the tool provides a rich platform for Java code clone experimentation including various configuration parameters like, e.g., used distance metrics, hashing algorithms, code representations.